OpenStudy (anonymous):

Fan & Medal / Please helpppp Find the circumference and area of a circle with a diameter of 22 inches. Leave your answers in terms of pi. C = 11π; A = 44π C = 22π; A = 44π C = 11π; A = 121π C = 22π; A = 121π

OpenStudy (anonymous):

@Cosmichaotic

OpenStudy (anonymous):

Circumfrence is 2pi(r) C = 2pi(r) Radius in diameter / 2 r = d/2 -> r = 22/2 -> r = 11 So C = 2pi(11)

OpenStudy (anonymous):

Which is C = 22pi Area is pi(r)^2 We know r = 11 A = pi(11)^2 A = pi(121)

OpenStudy (anonymous):

\[C = 22\pi \] \[A = 121\pi \]

OpenStudy (anonymous):

thank you so much!!!